Reciprocal Method of Support Department Cost Allocation

Valron Company has two support departments, Human Resources and General Factory, and two producing departments, Fabricating and Assembly.



Support Departments
Producing Departments

Human
Resources
General
Factory
FabricatingAssembly
Direct costs$170,000$330,000$114,200$95,000
Normal activity:



Number of employees6080170
Square footage1,0005,70013,300

The costs of the Human Resources Department are allocated on the basis of number of employees, and the costs of General Factory are allocated on the basis of square footage.

Now assume that Valron Company uses the reciprocal method to allocate support department costs.

Required:

1. Calculate the allocation ratios (rounded to four significant digits) for the four departments using the reciprocal method. If an amount is zero, enter "0". Use the rounded values for subsequent calculations.


Proportion of Driver Used by

Human ResourcesGeneral FactoryFabricatingAssembly
Human Resources



General Factory



2. Develop a simultaneous equations system of total costs for the support departments. If required, round your answers to four decimal places. Use these numbers for subsequent calculations. If required, round all other intermediate calculations to six decimal places, except the answers computed in requirement 1.

Human Resources (HR)=$+(GF)
General Factory (GF)=$+(HR)

Solve for the total reciprocated costs of each support department. (Round reciprocated total costs to the nearest dollar.)

Human Resources (HR)$
General Factory (GF)$

3. Using the reciprocal method, allocate the costs of the Human Resources and General Factory departments to the Fabricating and Assembly departments. Round all allocated costs to the nearest dollar. If an amount is zero, enter "0". Note: There may be a "$1" difference due to intermediate rounding.


Support Departments
Producing Departments

Human ResourcesGeneral Factory
FabricatingAssembly
Direct costs$$
$$
Allocate:




  Human Resources    
    
  General Factory    
    
Total after allocation$$
$$
Answer #1

1.

Proportion of Driver Used by
Human Resources General Factory Fabricating Assembly
Human Resources NIL 60/(60+80+170) = 0.194 80/310 = 0.258 170/310 = 0.548
General Factory 1000/(1000+5700+13300) = 0.05 NIL 5700/20000 = 0.285 13300/20000 = 0.665

2.

Human Resources (HR) = $ 170000 + 0.05 (GF)
General Factory (GF) = $ 330000 + 0.194 (HR)

Solving above, HR = 170,000 + 0.05 (330,000 + 0.194 HR)

=> HR = 186500 + 0.0097 HR

So HR = $ 188327 and GF = $ 366535.

3.


Support Departments
Producing Departments
Human
Resources
General
Factory
Fabricating Assembly
Direct costs $170,000 $330,000 $114,200 $95,000
Allocation
HR (188327) 188327*0.194 = 36535 188327*0.258 = 48589 188327*0.548=103203
GF 366535*0.05=18327 (366535) 366535*0.285 = 104462 366535*0.665 = 243746
Total NIL NIL 267251 441949